OpenStudy (anonymous):

Let sequence $(a_{n}): 0,1,1,2,2,3,3,4,4,....,k,k,k+1,k+1,.....$. Compute the sum $S_{n}=a_{1}+a_{2}+.....+a_{n}$.

OpenStudy (anonymous):

\[n(n+1)\times\frac{1}{2}\times 2\]

OpenStudy (asnaseer):

isn't this just twice the sum of a simple arithmetic progression: 1, 2, 3, 4, ..., n?

OpenStudy (zarkon):

if you were adding up from 1 to 2n

OpenStudy (asnaseer):

oh yes, so its twice the sum of: 1, 2, 3, ..., n/2

OpenStudy (zarkon):

it is thrown off a little since the first value is 0

OpenStudy (asnaseer):

but the zero seems to be ignored because it says S = a(1) + a(2) + ...

OpenStudy (zarkon):

is \[a_1\text{ or }a_0\] the first term?

OpenStudy (asnaseer):

\[a_1\]I see what you are saying now, so a1 IS zero

OpenStudy (zarkon):

that's what I think

OpenStudy (asnaseer):

so my guess is the sum will involve some mod operators to check for odd/even n?

OpenStudy (zarkon):

yes

OpenStudy (asnaseer):

ok, so taking @Zarkon's advice, here is my attempt at this: if n is odd, then there are (n+1)/2 terms starting with 0 and (n-1)/2 terms starting with 1. so we get the Odd N Sum (ONS) as:\[ONS_n=\frac{n+1}{4}(0+(\frac{n+1}{2}-1))+\frac{n-1}{4}(2+(\frac{n-1}{2}-1))\]\[=\frac{n^2-1}{8}\] if n is even, then there are n/2 terms starting with 0 and n/2 terms starting with 1. so we get the Even N Sum (ENS) as:\[ENS_n=\frac{n}{4}(0+(\frac{n}{2}-1))+\frac{n}{4}(2+(\frac{n}{2}-1))\]\[=\frac{n^2}{4}\] so: n MOD 2 == 0 ==> S_n = ESN_n = n^2/4 n MOD 2 != 0 ==> S_n = ONS_n = (n^2-1)/8

OpenStudy (asnaseer):

these all make use of the sum of an arithmetic progression to 'n' terms where the difference between each term is 'd' and the first term is 'a' being:\[S_n=\frac{n}{2}(2a+(n-1)d)\]in this case 'd' is 1 for both sequences and the first term 'a' is 0 for one sequence and 1 for the other.

OpenStudy (asnaseer):

sorry I wrote the sum for Odd 'n' incorrectly, it be (n^2-1)/4 I guess we could simplify it further by writing the sum as: S_n = (n^2 - (n MOD 2))/4 which takes care of the odd/even nature of 'n'.

OpenStudy (asnaseer):

does anyone know of how to enter MOD in the equation editor?

OpenStudy (asnaseer):

I managed to work out how to enter mod in Latex, so the sum can be simplified to:\[S_n=\frac{n^2-(n\mod{2})}{4}\]

OpenStudy (anonymous):

\[s_{n}=n^{2}/4 , or (n^{2}-1)/4\]

OpenStudy (asnaseer):

for even values of n, \(S_n=\frac{n^2}{4}\), and for odd values of n, \(S_n=\frac{n^2-1}{4}\). these can be combined into a single expression using the "mod" operator as I showed above.